# Jinhe Bi1,2,4  Danqi Yan1  Yifan Wang1  Aniri1  Wenke Huang3  Zengjie Jin1  Xiaowen Ma1 Artur Hecker2Mang Ye3   Xun Xiao2Hinrich Schütze1,4 Volker Tresp1,4   Yunpu Ma1,411footnotemark: 1

### v2

@@ Abstract
Jinhe Bi1,2,4   Danqi Yan1   Yifan Wang1   Aniri1   Wenke Huang3   Zengjie Jin1   Xiaowen Ma1 Artur Hecker2Mang Ye3   Xun Xiao2thanks: Corresponding authors: cognitive.yunpu@gmail.com, drxiaoxun@gmail.com.Hinrich Schütze1,4 Volker Tresp1,4   Yunpu Ma1,411footnotemark: 1 1 Ludwig Maximilian University of Munich  2 Munich Research Center, Huawei Technologies 3 School of Computer Science, Wuhan University  4 Munich Center for Machine Learning
@@ 1 Introduction
Figure 2: Visual Diagnosis of Representation Anisotropy in Pre-trained MLLM Features. (Left) The per-dimension mean distribution reveals a fundamental disparity: unlike text features (grey) which are well-centered around zero, all visual instruction datasets exhibit a significant non-zero mean. This provides direct evidence that visual embeddings occupy a biased, narrow cone, which is the geometric origin of Global Semantic Drift. (Right) The singular value scree plot confirms this diagnosis. The sharp "elbow point" indicates that the feature variance is confined to a few dominant dimensions—a classic symptom of representation degeneration.
The advent of MLLMs marks a significant milestone in artificial intelligence, endowing models with remarkable capabilities to process and reason over interleaved visual and textual information (Liu et al., 2024a; Zhu et al., 2023; Dai et al., 2023; Bi et al., 2025a; Wang et al., 2025). The standard paradigm for building these models is a two-stage process: (1) large-scale pre-training on web-scraped image-text pairs to establish foundational cross-modal alignment, followed by (2) visual instruction tuning to hone the model’s ability to follow complex, task-oriented instructions. While instruction tuning is critical for unlocking state-of-the-art performance, its effectiveness is increasingly challenged by the data it relies on. Publicly curated datasets are expanding exponentially but are often saturated with low-quality and redundant examples (Chen et al., 2024; Wei et al., 2023). Naively training on this noisy data inflates computational costs and, paradoxically, can degrade final model performance. This predicament has created an urgent need for efficient visual instruction selection methods that can identify a compact, high-quality subset of data.
@@
  • Our experiments show PRISM sets a new state-of-the-art, reducing training time by 70% while yielding performance superior to models fine-tuned on the complete dataset across numerous benchmarks.
@@ 2 Visual Instruction Selection
Figure 3: End-to-end computational cost (GPU Hours) vs. final model performance. PRISM achieves state-of-the-art performance while significantly reducing total time, thus satisfying the condition OSC < 1.
@@
As illustrated in Fig. 4, existing visual instruction selection methods can be broadly categorized into two main paradigms. The first is Proxy-Based Selection, where the target MLLM remains a black box, and data quality is evaluated using an external scoring function. This function can be a purpose-built scoring model (Chen et al., 2024), a human reward model, or a powerful LLM-based scorer (Wei et al., 2023). Some approaches also train a smaller proxy MLLM to guide the data selection for a larger target model (Lee et al., 2024). The second category is Training-Based Selection. In this approach, the target MLLM is actively used to assess data value by computing metrics derived from its training dynamics, such as training loss, perplexity, or gradient information on a given data partition. Yet both paradigms exhibit structural drawbacks: they risk performance degradation, incur high computational costs, and introduce Proxy-Based bias. A fuller discussion appears in Appendix D. We contend that these drawbacks share a single, underappreciated root cause: the overlooked representation anisotropy inherent in MLLM visual feature distributions. To address this geometric pathology, we introduce PRISM—a self-PRuning Intrinsic Selection Method—that enables training-free multimodal data selection by operating directly on the model’s intrinsic, re-centered features.
3 [Uncaptioned image]PRISM
@@ 3.2 Theoretical Analysis: Anisotropy as a Source of Computational Overhead
where $\boldsymbol{\delta}^{\prime}$ is the component of $\boldsymbol{\delta}$ orthogonal to $\boldsymbol{\mu}$. The true semantic similarity, proportional to $\boldsymbol{\delta}_{i}^{\top}\boldsymbol{\delta}_{j}$, is relegated to lower-order terms, rendering the metric ineffective for distinguishing unique samples. A full derivation is provided in Appendix E. ∎
@@ 4 Experiments
Model SQA SQA-I VizWiz POPE-P POPE-R POPE-A $M$-Vet MMBench MME-P MME-C MMMU Rel. (%)
LLaVA-Phi2-3B 75.3 72.7 41.2 87.3 88.6 86.1 35.6 68.7 1467.7 298.0 37.7 100%
[Uncaptioned image]PRISM-3B 76.3 72.8 40.9 87.5 88.8 86.5 34.1 68.9 1485.5 305.0 37.6 100.1%
LLaVA-Vicuna-7B 69.4 66.8 50.0 86.1 87.3 84.2 31.1 64.3 1510.7 311.9 35.4 100%
[Uncaptioned image]PRISM-7B 71.3 69.1 50.1 87.7 88.7 85.5 32.0 65.2 1470.0 330.0 34.7 101.7%
LLaVA-Vicuna-13B 74.4 71.6 53.6 87.4 88.0 85.6 36.1 67.7 1531.3 295.4 35.1 100%
[Uncaptioned image]PRISM-13B 74.5 71.8 53.1 87.7 88.4 85.7 36.4 65.8 1538.5 307.5 35.7 100.4%
Table 2: PRISM discovers a universal, model-agnostic data core. The PRISM-selected data for all experiments was generated only once using the LLaVA-1.5-7B model. This single, fixed subset consistently enhances performance not only on the source model but also when transferred to models of different scales (3B, 13B) and architectures (Phi-2). This validates that by correcting intrinsic feature properties, PRISM identifies data that is fundamentally valuable, transcending specific model configurations.
@@ 4.1 Experiment Setup
Benchmarks. Following the rigorous framework of LLaVA-1.5 (Liu et al., 2024a), we evaluate PRISM across a comprehensive suite of seven diverse benchmarks. This suite is designed to probe a wide array of MLLM capabilities, including fine-grained perception, visual reasoning, hallucination detection, and complex conversational skills. Our evaluation includes MMBench (Liu et al., 2024c), ScienceQA (Lu et al., 2022), MME (Yin et al., 2023), POPE (Li et al., 2023), VizWiz (Gurari et al., 2018), MM-Vet (Yu et al., 2024b), and MMMU (Yue et al., 2024). Detailed descriptions of each benchmark are provided in Appendix F.
4.2 Main Results
Balanced Efficiency and Performance. As illustrated in Figure 3, PRISM strikes an optimal balance between computational cost and model performance. While strong Training-Based methods like TIVE can achieve comparable performance (100.6% rel.), their substantial selection overhead often causes the total end-to-end time to exceed that of full fine-tuning, violating our Principle of Net Efficiency Gain. In stark contrast, PRISM not only achieves superior accuracy (101.7% rel.) but also reduces the total pipeline time by approximately 70%. This exceptional efficiency is inherent to our training-free design, where feature extraction and correlation computations are performed in a single offline pass with negligible overhead compared to iterative training. A detailed breakdown of the wall-clock runtime can be found in Appendix G.
Model Hellaswag MMLU MMLU-PRO Rel. (%)
LLaVA-Phi2-3B 66.0 50.5 9.1 100%
[Uncaptioned image]PRISM-3B 67.4 52.7 8.6 100.3%
LLaVA-Vicuna-7B 66.5 35.0 17.8 100%
[Uncaptioned image]PRISM-7B 66.5 41.1 15.7 101.9%
LLaVA-Vicuna-13B 69.5 36.2 6.8 100%
[Uncaptioned image]PRISM-13B 69.6 39.5 12.4 130.6%
4.3 In-Depth Model Behavior Analysis
Cross-Model Generalization and Scalability. A key strength of PRISM is its ability to identify a universally high-value data subset, rather than one overfitted to a specific model. This stems from our method’s core principle: by correcting for a fundamental geometric flaw (anisotropy) and operating on intrinsic feature correlations, PRISM isolates data with inherent semantic value that transcends architectural specifics. To validate this, we apply the subset selected using LLaVA-1.5-7B to different model architectures and scales (details in Appendix 6). As shown in Table 2, the performance gains remain robust, confirming that the PRISM-selected data is highly transferable and scalable.
@@ 4.5 Ablation Study
Impact of Correlation-based Selection. Validating our central hypothesis, we show that selecting the Low-Correlation data subset yields the highest performance. This confirms that our implicit re-centering strategy effectively identifies informational diversity and prunes redundant samples that provide diminishing returns for model training (see Appendix D for further analysis).
Effect of Token Aggregation Strategy. We confirm that using an Average of all image tokens to form the global feature vector is superior to using only the Last Image Token. Averaging provides a more holistic and robust representation of an image’s semantics, which is crucial for the stability of our geometric analysis, whereas relying on a single token can be susceptible to positional biases (Bi et al., 2025b).
@@ 5 Conclusion
We identify that inefficient data selection for MLLMs is rooted in a fundamental geometric flaw: representation anisotropy. We propose PRISM, a training-free, intrinsic selection method that corrects this flaw via implicit re-centering to reveal the data’s underlying semantic structure. Empirically, PRISM reduces training time by over 70% while yielding models that outperform those trained on the full, unpruned dataset. Our work thus advocates for a paradigm shift from complex selection heuristics towards more fundamental, geometry-aware data curation.
@@ Appendix A Related Work
Visual Instruction Selection: Despite the strong performance of MLLMs, the rapid growth of visual instruction datasets has introduced significant redundancy, similar to challenges in LLMs (Zhou et al., 2024; Xia et al., 2024). State-of-the-art models like BLIP3 (Xue et al., 2024), InternVL2.5 (Chen et al., 2025), and LLaVA-OneVision (Li et al., 2024) rely on billions of instructions to enhance understanding, but their massive scale leads to substantial computational costs, often requiring hundreds to thousands of GPU hours.
@@ B.2 Final Dataset Composition
Appendix C Model Architecture Details
To rigorously evaluate the generalization capabilities of PRISM, we selected a diverse set of three open-source Multimodal Large Language Models (MLLMs). The goal was to test whether the data subset curated by PRISM remains effective when transferred across different model scales, language model families, and vision encoder architectures. This diversity ensures that our method’s effectiveness is not tied to a single model configuration but rather to the fundamental data properties it identifies.
Our selection, summarized in Table 6, includes models from the LLaVA family with varying components:
  • Language Model Scale: We test on 3B, 7B, and 13B parameter language models to assess scalability.
  • Language Model Family: We use models from two distinct families, Vicuna (Chiang et al., 2023) (based on Llama) and Phi-2 (Valipour et al., 2023), to validate architectural transferability.
  • Vision Encoder: We include models built on different vision backbones, specifically CLIP (Radford et al., 2021) and the more recent SigLIP (Zhai et al., 2023), to ensure PRISM is robust to the choice of vision foundation model.
Table 6: Architectural configurations of MLLMs for evaluating generalization. Our selection spans different language model families (Vicuna vs. Phi-2), scales (3B, 7B, 13B), and vision encoders (CLIP vs. SigLIP) to ensure a robust validation.
Model Vision Encoder Language Model
LLaVA-Vicuna-7B CLIP ViT-L/14 336px (Radford et al., 2021) Vicuna v1.5 7B (Chiang et al., 2023)
LLaVA-Phi2-3B SigLIP-SO400M-Patch14-384 (Zhai et al., 2023) Phi-2 2.7B (Valipour et al., 2023)
LLaVA-Vicuna-13B CLIP ViT-L/14 336px (Radford et al., 2021) Vicuna v1.5 13B (Chiang et al., 2023)
Appendix D Limitations of Existing Methods and Our Geometric Motivation

### v3

# Jinhe Bi1,2,4  Aniri1  Danqi Yan1  Yifan Wang1  Wenke Huang3 Zengjie Jin1Xiaowen Ma1Sikuan Yan1Artur Hecker2Mang Ye3 Xun Xiao2Hinrich Schütze1,4 Volker Tresp1,4   Yunpu Ma1,411footnotemark: 1

@@ Abstract
Jinhe Bi1,2,4   Aniri1   Danqi Yan1   Yifan Wang1   Wenke Huang3 Zengjie Jin1Xiaowen Ma1Sikuan Yan1Artur Hecker2Mang Ye3 Xun Xiao2thanks: Corresponding authors: cognitive.yunpu@gmail.com, drxiaoxun@gmail.com.Hinrich Schütze1,4 Volker Tresp1,4   Yunpu Ma1,411footnotemark: 1 1 Ludwig Maximilian University of Munich  2 Munich Research Center, Huawei Technologies 3 School of Computer Science, Wuhan University  4 Munich Center for Machine Learning
@@ 1 Introduction
Figure 2: Visual Diagnosis of Representation Anisotropy in Pre-trained MLLM Features. (Left) The per-dimension mean distribution reveals a fundamental disparity: unlike text features (grey) which are well-centered around zero, all visual instruction datasets exhibit a significant non-zero mean. This provides direct evidence that visual embeddings occupy a biased, narrow cone, which is the geometric origin of Global Semantic Drift. (Right) The singular value scree plot confirms this diagnosis. The sharp "elbow point" indicates that the feature variance is confined to a few dominant dimensions—a classic symptom of representation degeneration. Detailed statistics and qualitative case studies are provided in Appendix G and Fig. 8.
The advent of MLLMs marks a significant milestone in artificial intelligence, endowing models with remarkable capabilities to process and reason over interleaved visual and textual information (Liu et al., 2024a; Zhu et al., 2023; Dai et al., 2023). The standard paradigm for building these models is a two-stage process: (1) large-scale pre-training on web-scraped image-text pairs to establish foundational cross-modal alignment, followed by (2) visual instruction tuning to hone the model’s ability to follow complex, task-oriented instructions. While instruction tuning is critical for unlocking state-of-the-art performance, its effectiveness is increasingly challenged by the data it relies on. Publicly curated datasets are expanding exponentially but are often saturated with low-quality and redundant examples (Chen et al., 2024; Wei et al., 2023). Naively training on this noisy data inflates computational costs and, paradoxically, can degrade final model performance. This predicament has created an urgent need for efficient visual instruction selection methods that can identify a compact, high-quality subset of data.
@@
  • Our experiments show PRISM sets a new state-of-the-art, reducing training time by 70% while yielding performance superior to models fine-tuned on the complete dataset.
@@ 2 Visual Instruction Selection
Figure 3: End-to-end computational cost (GPU Hours) vs. final model performance. PRISM achieves state-of-the-art performance while significantly reducing total time, thus satisfying the condition OSC <1<1.
@@
As illustrated in Fig. 4, existing visual instruction selection methods can be broadly categorized into two main paradigms. The first is Proxy-Based Selection, where the target MLLM remains a black box, and data quality is evaluated using an external scoring function. This function can be a purpose-built scoring model (Chen et al., 2024), a human reward model, or a powerful LLM-based scorer (Wei et al., 2023). Some approaches also train a smaller proxy MLLM to guide the data selection for a larger target model (Lee et al., 2024). The second category is Training-Based Selection. In this approach, the target MLLM is actively used to assess data value by computing metrics derived from its training dynamics, such as training loss, perplexity, or gradient information on a given data partition. Yet both paradigms exhibit structural drawbacks: they risk performance degradation, incur high computational costs, and introduce Proxy-Based bias. A fuller discussion appears in Appendix E. We contend that these drawbacks share a single, underappreciated root cause: the overlooked representation anisotropy inherent in MLLM visual feature distributions. To address this geometric pathology, we introduce PRISM—a self-PRuning Intrinsic Selection Method—that enables training-free multimodal data selection by operating directly on the model’s intrinsic, re-centered features.
3 PRISM
@@ 3.2 Theoretical Analysis: Anisotropy as a Source of Computational Overhead
where $\boldsymbol{\delta}^{\prime}$ is the component of $\boldsymbol{\delta}$ orthogonal to $\boldsymbol{\mu}$. The true semantic similarity, proportional to $\boldsymbol{\delta}_{i}^{\top}\boldsymbol{\delta}_{j}$, is relegated to lower-order terms, rendering the metric ineffective for distinguishing unique samples. A full derivation is provided in Appendix F. ∎
@@ 4 Experiments
Figure 5: Generalization across diverse MLLM families and scales. Comparison of PRISM (blue) and full fine-tuning (grey) across SQA-I, POPE, MMBench, and Rel. (%) (detailed numerical results provided in Table 6 ). Results for Phi-2, Vicuna, Llama-3, and Qwen2.5 confirm that PRISM’s performance gains are architecture-agnostic and transfer robustly across varied backbones and vision encoders.
@@ 4.1 Experiment Setup
Benchmarks. Following the rigorous framework of LLaVA-1.5 (Liu et al., 2024a), we evaluate PRISM across a comprehensive suite of seven diverse benchmarks. This suite is designed to probe a wide array of MLLM capabilities, including fine-grained perception, visual reasoning, hallucination detection, and complex conversational skills. Our evaluation includes MMBench (Liu et al., 2024c), ScienceQA (Lu et al., 2022), MME (Yin et al., 2023), POPE (Li et al., 2023), VizWiz (Gurari et al., 2018), MM-Vet (Yu et al., 2024b), and MMMU (Yue et al., 2024). Detailed descriptions of each benchmark are provided in Appendix H.
4.2 Main Results
Balanced Efficiency and Performance. As illustrated in Figure 3, PRISM strikes an optimal balance between computational cost and model performance. While strong Training-Based methods like TIVE can achieve comparable performance (100.6% rel.), their substantial selection overhead often causes the total end-to-end time to exceed that of full fine-tuning, violating our Principle of Net Efficiency Gain. In stark contrast, PRISM not only achieves superior accuracy (101.7% rel.) but also reduces the total pipeline time by approximately 70%. This exceptional efficiency is inherent to our training-free design, where feature extraction and correlation computations are performed in a single offline pass with negligible overhead compared to iterative training. A detailed breakdown of the wall-clock runtime can be found in Appendix K.
Model Hellaswag MMLU MMLU-PRO Rel. (%)
LLaVA-Phi2-3B 66.0 50.5 9.1 100%
[Uncaptioned image]PRISM-3B 67.4 52.7 8.6 100.3%
LLaVA-Vicuna-7B 66.5 35.0 17.8 100%
[Uncaptioned image]PRISM-7B 66.5 41.1 15.7 101.9%
LLaVA-Vicuna-13B 69.5 36.2 6.8 100%
[Uncaptioned image]PRISM-13B 69.6 39.5 12.4 130.6%
4.3 In-Depth Model Behavior Analysis
Cross-Model Generalization and Scalability. PRISM’s selection mechanism generalizes robustly across diverse MLLM architectures by correcting representation anisotropy. By identifying semantically dense samples, we establish the approach’s breadth across five model–encoder combinations. Results demonstrate that effectiveness holds for both small and large language models—including Vicuna, Phi-2, Qwen2.5, and Llama-3—and for distinct vision encoders like CLIP and SigLIP (details in Appendix 5). These findings, illustrated in Fig.5, confirm that PRISM’s gains are not specific to the LLaVA architecture but transfer effectively across diverse multimodal families.
@@ 4.5 Ablation Study
Impact of Correlation-based Selection. Validating our central hypothesis, we show that selecting the Low-Correlation data subset yields the highest performance. This confirms that our implicit re-centering strategy effectively identifies informational diversity and prunes redundant samples that provide diminishing returns for model training (see Appendix E for further analysis).
Effect of Token Aggregation Strategy. We confirm that using an Average of all image tokens to form the global feature vector is superior to using only the Last Image Token. Averaging provides a more holistic and robust representation of an image’s semantics, which is crucial for the stability of our geometric analysis, whereas relying on a single token can be susceptible to positional biases.
@@ 5 Conclusion
We trace MLLM selection inefficiency to representation anisotropy and propose PRISM, a training-free method that restores semantic structure via implicit re-centering. Empirically, PRISM reduces training time by 70% while surpassing full-dataset performance , advocating for a shift toward geometry-aware curation over complex heuristics.
@@ Appendix A Related Work
Visual Instruction Selection: Despite the strong performance of MLLMs, the rapid growth of visual instruction datasets has introduced significant redundancy, similar to challenges in LLMs (Zhou et al., 2024; Chen et al., 2023; Xia et al., 2024; Tian et al., 2025; Lu et al., 2025; Bi et al., 2025b). State-of-the-art models like BLIP3 (Xue et al., 2024), InternVL2.5 (Chen et al., 2025c), and LLaVA-OneVision (Li et al., 2024) rely on billions of instructions to enhance understanding, but their massive scale leads to substantial computational costs, often requiring hundreds to thousands of GPU hours.
@@ B.1 Redundancy Score Distribution
Top-Ranked Sample
[Question] What region of the world is this located?
[Answer] Asia
[Reasoning] Requires fine-grained OCR and cultural reasoning.
Bottom-Ranked Sample
[Question] What type of weather is this?
[Answer] Sunny
[Reasoning] Trivial perception; high semantic redundancy.
Figure 8: Representative examples selected by PRISM from LLaVA-665K. Top-ranked samples correspond to more challenging and informative interactions that exhibit unique semantic structures, whereas bottom-ranked samples tend to be trivial, binary, or semantically redundant.
@@ B.2 Final Dataset Composition
B.3 Qualitative Analysis of PRISM Selection
To provide a more intuitive understanding of the Intrinsic Selection paradigm, we conduct a qualitative analysis of samples ranked by PRISM. Figure 8 illustrates representative examples from the top-ranked (selected) and bottom-ranked (pruned) tiers of the LLaVA-665K dataset.
B.3.1 High-Value: Challenging and Informative
The top-ranked samples prioritized by PRISM typically involve fine-grained perception and complex contextual reasoning. For instance, in the "Asia street" example (Figure 8, Left), the model must identify localized visual cues such as the road marking Tomare, the specific structure of utility poles, and architectural styles to correctly determine the geographical region. From a training perspective, these informative interactions provide a stronger learning signal, often corresponding to higher initial loss values, which is critical for mastering expert-level reasoning.
B.3.2 Low-Value: Trivial and Redundant
Conversely, bottom-ranked samples are frequently characterized by trivial perception or high semantic redundancy. In the "Sunny weather" example (Figure 8, Right), the task involves identifying a generic weather condition in a common park setting. Such visual scenes—comprising grass, trees, and sunlight—are over-represented in the corpus, leading to a high correlation with the Global Semantic Drift.
Because these samples offer saturated information and redundant gradients, training on them yields diminishing returns and may even exacerbate "knowledge forgetting". By pruning these trivial samples, PRISM ensures a more efficient use of the computational budget without sacrificing model performance.
Appendix C Model Architecture Details and Generalization Setup
To rigorously evaluate the generalization capabilities of PRISM, we extend our analysis across a diverse set of Multimodal Large Language Models (MLLMs). The objective is to verify whether the visual anisotropy diagnosis and the subsequent performance gains from PRISM remain consistent across different model scales, language model families, and vision encoder architectures. This broad evaluation ensures that the intrinsic semantic properties identified by PRISM are not artifacts of a specific model configuration but represent fundamental characteristics of multimodal data representations.
As summarized in Table 5, our evaluation framework encompasses the following dimensions:
  • Language Model Families and Scales: Beyond the Vicuna and Phi-2 families discussed in the main text, we include the Qwen2.5 series (7B Base and Instruct) and Llama-3 (8B) to represent state-of-the-art open-source LLM backbones.
  • Vision Encoders: We evaluate models utilizing various vision foundation backbones, specifically CLIP ViT-L/14 (336px) and SigLIP-SO400M, to ensure PRISM’s robustness to different cross-modal alignment strategies.
  • Generalization Analysis: For each configuration, we conduct both a systematic anisotropy diagnosis (as detailed in Section 3.2) and an end-to-end evaluation of the performance improvements yielded by PRISM-selected data.
Model Family Vision Encoder Language Model Backbone
LLaVA-Vicuna-7B CLIP ViT-L/14 (336px) Vicuna v1.5 7B
LLaVA-Phi2-3B SigLIP-SO400M Phi-2 2.7B
LLaVA-Vicuna-13B CLIP ViT-L/14 (336px) Vicuna v1.5 13B
LLaVA-Qwen-7B SigLIP-SO400M Qwen2.5-7B (Base/Instruct)
LLaVA-Llama-8B CLIP ViT-L/14 (336px) Llama-3-8B
Table 5: Comprehensive architectural configurations for generalization testing. The evaluation spans multiple LLM families, varying parameter scales, and diverse vision encoders to validate the universal applicability of PRISM.
Appendix D Impact of Textual Representation and Joint Multimodal Selection
To address potential concerns regarding the role of textual prompts in data redundancy, this section provides: (1) a comparative analysis of visual vs. textual anisotropy, (2) the motivation for our visual-centric approach, and (3) additional experimental results explicitly incorporating textual features during selection.
D.1 Contrastive Anisotropy: Visual vs. Textual Modalities
Our diagnosis of representation anisotropy reveals a significant disparity between modalities. As illustrated in Figure 2, while visual features across all instruction datasets exhibit a non-zero mean vector $\mu$ with substantial magnitude, text features produced by the same MLLM are comparatively well-centered around the origin. Furthermore, the singular value spectrum of the textual embedding space is considerably flatter than that of the visual space.
Because these two modalities exhibit fundamentally different geometric behaviors, applying a single global re-centering operation to concatenated multimodal features would be suboptimal. A joint mean vector would be dominated by the severe visual drift, potentially over-correcting or distorting the near-isotropic textual geometry, thereby degrading the quality of similarity estimates. Consequently, PRISM focuses its geometric correction on the visual stream, where anisotropy is the primary obstacle to redundancy detection.
D.2 Primacy of Visual Redundancy in Instruction Tuning
In the context of visual instruction tuning, the image typically serves as the primary carrier of new scene-level semantics. In datasets such as LLaVA-665K, text prompts often function as task-specific templates (e.g., questions or instructions) conditioned on the visual content. Therefore, redundancy in these datasets is largely governed by the visual channel: models frequently encounter near-identical visual semantics under slight variations of textual phrasing. By targeting visual anisotropy, PRISM effectively prunes redundant visual scenes that provide diminishing returns during the fine-tuning process.
D.3 Additional Experiments with Multimodal Features
To empirically verify this design choice, we conducted additional experiments on LLaVA-1.5-7B using features that explicitly incorporate both visual and textual information. We constructed multimodal representations by aggregating image tokens and text prompt tokens before applying the PRISM selection pipeline.
As shown in Table 7, including textual prompts in the selection signal does not yield performance gains over our visual-only approach. In fact, the multimodal-feature selection (97.8% relative performance) underperforms the visual-only PRISM (101.7%). We hypothesize this is because text prompts introduce additional variance related to syntactic style and template length that is uncorrelated with the underlying semantic redundancy of the task. These results reinforce that visual anisotropy is the dominant factor to address for training-free intrinsic selection.
Model SQA SQA-I VizWiz POPE-P POPE-R POPE-A $M$-Vet MMBench MME MMMU Rel. (%)
LLaVA-Phi2-3B 75.3 72.7 41.2 87.3 88.6 86.1 35.6 68.7 1765.7 37.7 100%
[Uncaptioned image]PRISM-3B 76.3 72.8 40.9 87.5 88.8 86.5 34.1 68.9 1790.5 37.6 100.1%
LLaVA-Vicuna-7B 69.4 66.8 50.0 86.1 87.3 84.2 31.1 64.3 1822.6 35.4 100%
[Uncaptioned image]PRISM-7B 71.3 69.1 50.1 87.7 88.7 85.5 32.0 65.2 1800.0 34.7 101.7%
LLaVA-Vicuna-13B 74.4 71.6 53.6 87.4 88.0 85.6 36.1 67.7 1826.7 35.1 100%
[Uncaptioned image]PRISM-13B 74.5 71.8 53.1 87.7 88.4 85.7 36.4 65.8 1846.0 35.7 100.4%
LLaVA-Qwen-7B (Base) 79.1 76.7 55.3 87.3 89.1 86.4 33.1 70.1 1923.1 46.0 100%
[Uncaptioned image]PRISM-Qwen-7B (Base) 81.2 78.9 55.1 88.6 90.3 87.5 33.7 71.5 1895.8 45.2 101.0%
LLaVA-Qwen-7B (Instruct) 80.2 77.8 54.7 87.9 89.7 86.9 33.8 71.0 1935.4 46.8 100%
[Uncaptioned image]PRISM-Qwen-7B (Instruct) 82.3 80.1 54.2 89.1 90.8 88.1 34.2 72.4 1908.6 46.1 100.9%
LLaVA-Llama-8B 77.6 75.2 53.7 86.3 87.6 84.0 31.8 68.5 1889.7 40.3 100%
[Uncaptioned image]PRISM-Llama-8B 79.4 77.3 53.4 87.8 89.1 85.6 31.5 69.9 1862.5 39.7 100.8%
Table 6: Cross-Model Generalization and Scalability of PRISM. PRISM’s selection mechanism generalizes robustly across diverse MLLM architectures by correcting representation anisotropy, as shown across five LLaVA-style model–encoder combinations. By identifying semantically dense samples for each architecture, PRISM consistently improves performance over full finetuning for both small and large language models—including Vicuna, Phi-2, Qwen2.5, and Llama-3—as well as distinct vision encoders such as CLIP and SigLIP. Together with Fig. 5, these results confirm that PRISM’s gains are not tied to a particular LLaVA configuration but transfer across diverse multimodal model families.
Multimodal PRISM 68.2 65.1 48.9 85.3 86.0 83.1 30.2 63.0 1455.0 305.0 34.1 97.8%
[Uncaptioned image] Visual-only PRISM 71.3 69.1 50.1 87.7 88.7 85.5 32.0 65.2 1470.0 330.0 34.7 101.7%
Table 7: Comparison of selection using visual-only features versus joint multimodal features. The visual-centric approach (PRISM) demonstrates higher stability as text prompts often introduce non-semantic noise.
Appendix E Limitations of Existing Methods and Our Geometric Motivation